Matthew Gray

38 8th Street, Apt 7 Cambridge, MA 02141

October 3, 1996

Sir or Madam:

Thank you for taking the time to read this resume. With the range of technical and business experience I have, I think you will find I am a skilled, motivated and responsible individual who would be a valuable asset to your company.

As you can see from my resume, I have a range of work experiences, focusing on physics and software development, but I have also been involved in nearly every aspect of the business process through my own company, net.Genesis which I started in 1994 (which still does business, but I have since left to complete my education). I have learned to extend my technical skills and expertise to communicate within a business organization, particularly to sales and marketing. I have also gained experience in managing teams of engineers, handling investor relationships and interactions with the press. Additionally, as the world’s leading authority on growth of the web (see Publications), I have had the opportunity to speak at conferences, coordinate with others in the technical community and learn the value of good PR.

I have leadership abilities that have developed through experiences including net.Genesis and my chairmanship of the SIPB. Additionally, I think you will find my traditional technical skills exceptional, due to a lifetime of software experience and an extensive expermintal and theoretical research background.

I hope you will find my technical skills outstanding and my industry experience valuable and relevant. Thank you once again for taking the time to read this, and I look forward to hearing from you so that we may schedule a personal interview.

Sincerely,

Matthew Gray